# Exploiting the Semantic Knowledge of Pre-trained Text-Encoders for Continual Learning

### v1

@@ 5.1 Experimental Settings
TABLE IV: Results on eight fine-grained datasets with different task splits.
@@
Dataset Food Cars Aircraft Pets Caltech Flowers CUB Dogs
#classes 101 196 100 37 102 102 200 120
Linear probe 92.8 86.7 59.5 93.1 94.7 98.1 80.4 78.6
Joint 93.1 91.5 73.2 94.7 93.8 98.1 85.0 80.6
Continual-CLIP[22] 89.2 65.6 27.1 88.9 89.3 70.4 55.6 63.4
 CoOp[23] 84.6 86.1 57.6 94.0 90.2 97.0 80.1 75.8
CoCoOp[38] 79.1 84.9 59.8 92.8 88.2 97.2 77.0 69.7
L2P[14] 81.7 86.9 61.1 92.8 88.6 97.0 82.6 75.9
DualPrompt[15] 80.5 82.7 45.5 90.6 86.4 96.7 79.6 70.3
Ours 90.9 ($\uparrow$1.7) 88.6 ($\uparrow$2.5) 66.6 ($\uparrow$5.5) 94.0 92.3 ($\uparrow$2.1) 96.2 83.7 ($\uparrow$1.1) 79.6 ($\uparrow$3.7)
@@ 5.2 Full-shot Continual Learning Setting
We conducted extensive experiments on three benchmark datasets, namely CIFAR100, imagenet_subset, and ImageNet1000 in comparison to state-of-the-art approaches. In Table II, we present the averaged accuracy on CIFAR100 over three different class orders and three different splits. In the most common 10-step split, our method achieves an impressive accuracy of 80.1 after the last incremental step, surpassing the previous state-of-the-art method by 11.4 points. Even for longer task sequences of 20 or 50 steps, our method consistently maintains superior performance, with at least 9.5 and 5.4 points higher accuracy than the state-of-the-art ‘Last’ result respectively. Moving to larger-scale continual learning scenarios, as shown in Table III, our method demonstrates even greater superiority. On imagenet_subset dataset, we outperform the state-of-the-art result by 7.9 points in terms of top-1 ‘Last’ accuracy, approaching the performance of linear probing [19]. Furthermore, on the challenging ImageNet1000 dataset, our method achieved significant improvements with a 6.5-point increase in top-1 ‘Last’ accuracy and a 3.1-point increase in top-5 accuracy. Overall, our experimental results demonstrate that the proposed method consistently outperforms state-of-the-art approaches on all three datasets, showcasing its superior performance and stability in continual learning settings.
@@
TABLE V: The effect of different components. ‘FT’ represents only applying contrastive loss for training, ‘One-hot Label’ utilizes the prompt of each class as a classification head during training, and uses the one-hot encoded labels to calculate the cross-entropy loss. ‘Naive KD’ distills the output of the current data between old and new models with the KD loss proposed in[54] (ref Eq. 11).
@@
To further validate the effectiveness of the proposed method, we conduct experiments on eight fine-grained datasets. In fine-grained datasets, all categories belong to the same parent category in a hierarchical relationship, which intuitively indicates a high degree of semantic similarity among them, which is more challenging. In Table IV, we present the performance of our proposed method compared to other methods on different splits of the datasets. We notice that the Continual-CLIP evaluation on datasets like Cars, Aircrafts, and CUB, is initially poor. State-of-the-art methods obtain a substantial improvement, particularly on Aircraft where the gain reaches 34 points with L2P. Our method achieves performance very close to the joint training results, with a gap of less than 1 point on multiple datasets such as Pets and Dogs. Moreover, our method even surpasses the results of linear probing evaluation on multiple datasets, which is typically used to assess the quality of pre-trained features.
@@ 5.4.1 Components of our proposed method
We conduct ablation experiments on CIFAR100 and imagenet_subset to thoroughly examine the effect of the different proposed modules. Both datasets were split into 10 steps, and the results are shown in Table V. On both datasets, the contrastive loss significantly improves the final accuracy compared to Continual-CLIP. However, when we additionally apply the one-hot encoded labels to add an auxiliary classification loss, we observe a significant performance drop. By incorporating our ‘SG-RL’ module, the model shows a gain of 1.3 points on CIFAR100 and 0.7 points on imagenet_subset. Furthermore, when we apply knowledge distillation to the loss function, our proposed ‘SG-KD’ demonstrates superior performance over ‘Naive KD’ (ref Eq. 11) on both datasets, with a particularly notable gain of 2.9 points on Cifar100.
TABLE VI: Comparison of different vision models on CIFAR100 and imagenet_subset, learned with 10 steps.
CIFAR100 imagenet_subset
Avg Last Avg Last
ResNet 50 C-CLIP[22] - 41.6 - 65.6
Ours 59.8 50.7 80.2 71.3
ViT-B/16 C-CLIP[22] - 68.7 - 75.2
Ours 86.6 80.1 89.8 83.1
ViT-L/14 C-CLIP[22] - 77.9 - 81.5
Ours 91.1 86.5 91.9 85.8
TABLE VII: Comparison of training different layers of image encoder.
@@
Figure 5: The comparison of accuracy when saving different number of exemplars on CIFAR100 and imagenet_subset.
@@ 5.4.2 Validation on various vision models
Our primary experiments employed ViT-B/16 as the image encoder. However, it is important to investigate the effectiveness of our method with other architectures as well. Hence, we validate our approach on two additional vision models: ResNet 50 and ViT-L/14. Table VI presents the comparison between Continual-CLIP (illustrated as ‘C-CLIP’ for short in the table) evaluation and our proposed method on CIFAR100 and imagenet_subset within 10 steps. Despite the differences in architecture and parameter count, all three image encoders exhibit significant improvements compared to Continual-CLIP evaluation. These results suggest that our method is effective across different scales of image encoders, irrespective of architecture.
TABLE VIII: Ablation study on hyper-parameters sensitivity. (‘C’ stands for ‘CIFAR100’ and ‘I’ stands for ‘imagenet_subset’.)
@@
Dataset CIFAR100 imagenet_subset ImageNet1000 Food Cars Aircraft Pets Caltech Flowers CUB Dogs
#classes 100 100 1000 101 196 100 37 102 102 200 120
Linear probe [^5] We directly add a fully connected layer after the linear projection, we train 100 epochs for each dataset with an initial learning rate of 0.1 and decays by 0.1 every 45 epochs. 83.1 83.9 80.2 92.8 86.7 59.5 93.1 94.7 98.1 80.4 78.6
Joint 83.9 86.1 81.1 93.1 91.5 73.2 94.7 93.8 98.1 85.0 80.6
Continual-CLIP[22] 68.7 75.2 68.6 89.2 65.6 27.1 88.9 89.3 70.4 55.6 63.4
CoOp[23] 66.6 72.0 67.5 84.7 81.7 48.4 93.3 86.9 94.2 78.1 74.0
CoCoOp[38] 64.1 69.1 58.9 86.0 74.9 39.4 93.2 86.6 93.0 70.0 69.7
L2P[14] 57.8 59.7 58.8 74.9 63.6 27.3 83.3 73.1 77.6 70.0 58.9
DualPrompt[15] 67.4 71.3 65.6 85.2 80.9 49.9 90.5 90.5 96.8 76.2 69.8
Ours 80.6 83.1 75.1 90.9 88.6 66.6 94.0 92.3 96.2 83.7 79.6
TABLE IX: Results with frozen backbone on all datasets.
@@ 5.4.3 Effect of training different layers of image encoder
From Table VII, we observe that when training the last 6 blocks of the image encoder (half of the blocks in ViT-B/16), the model fails to overcome catastrophic forgetting. However, significant improvements are achieved when our proposed SG-RL and SG-KD modules are applied. Training only the linear projection layer yields decent results, but it appears that the model’s potential is greatly limited with such a small number of parameters. To explore the impact of incorporating more parameters into the training process, we gradually training additional blocks. It is evident that the best performance is achieved when only the last blocks are trained.
@@ 5.4.5 Hyper-parameters sensitivity
We conduct sensitivity analysis experiments on the two proposed losses to assess their impact. It can be seen from Table VIII, our method shows strong robustness to both hyper-parameters ranged from 0.1 to 1. The best results are obtained when $\lambda_{1}$ and $\lambda_{2}$ are set to 0.5, 0.1 respectively.
@@
5.5 More Analysis
5.5.1 Results with frozen backbone
Given that the methods we compared are all proposed based on frozen backbone while our method proposes to fine-tune the last block of the vision encoder. For a fair comparison, in the above experiment section we fine-tuned the same part of parameters for these methods. Here for a more comprehensive and in-depth comparison, we present the experimental results of these methods in the case of frozen backbone in Table IX. It can be clearly seen that adapting the backbone outperforms the frozen backbone by a large margin in vast majority of cases. It is particularly worth mentioning that on Out-Of-Distribution tasks such as Aircrafts and Cars, most methods show poor performance in the case of frozen backbone, indicating that these methods struggle to adapt to these tasks with only prompt steering pre-trained representations.
5.5.2 Comparison with one-hot label
@@
Dataset CIFAR100 imagenet_subset ImageNet1000
split $10\times 10$ $10\times 10$ $100\times 10$
One-hot Label 71.2 73.2 64.3
Ours 80.6 83.1 75.1
TABLE XI: Results compared to one-hot label on three general datasets with different task splits.

### v2

@@ 5.1.3 Evaluation metrics
TABLE IV: Last accuracy on eight fine-grained datasets with different task splits.
@@
Dataset ZS LB Food Cars Aircraft Pets Caltech Flowers CUB Dogs
#classes - - 101 196 100 37 102 102 200 120
split - - 11+$9\times 10$ $28\times 7$ $10\times 10$ 12+25 $17\times 6$ $17\times 6$ $20\times 10$ $12\times 10$
Linear probe - - 92.8 86.7 59.5 93.1 94.7 98.1 80.4 78.6
Joint - - 93.1 91.5 73.2 94.7 93.8 98.1 85.0 80.6
L2P [14] 74.9 63.6 27.3 83.3 73.1 77.6 70.0 58.9
DualPrompt [15] 85.2 80.9 49.9 90.5 90.5 96.8 76.2 69.8
RanPAC [46] 92.0 86.3 62.3 91.7 94.6 97.7 82.7 77.0
CODA-Prompt [23] 83.1 72.2 27.7 90.1 85.6 88.3 63.6 58.4
ConvPrompt [24] 84.1 86.3 53.7 91.1 88.4 95.4 79.1 71.1
EASE [25] 86.5 76.7 41.3 92.6 91.8 93.5 73.0 69.8
[…]
CoOp [55] 84.7 81.7 48.4 93.3 86.9 94.2 78.1 74.0
CoCoOp [54] 86.0 74.9 39.4 93.2 86.6 93.0 70.0 69.7
SPU [50] 77.2 68.5 59.7 93.1 69.6 90.0 66.3 67.4
CLAP4CLIP [29] 74.6 72.4 45.4 91.4 91.3 93.7 82.0 61.8
RAPF [28] 83.0 44.1 31.0 79.0 66.8 70.2 27.0 47.8
PROOF [51] 84.9 80.0 51.0 89.7 90.4 93.8 74.1 69.9
Ours 90.9 88.6 66.6 94.0 92.3 96.2 83.7 79.6
@@ 5.2 Full-shot Continual Learning Setting
We conducted extensive experiments on three benchmark datasets, namely CIFAR100, ImageNet_subset, and ImageNet1000 in comparison to state-of-the-art approaches. In Table II, we present the averaged accuracy on CIFAR100 over three different class orders and three different splits. We categorize all compared methods into three groups: (1) training from scratch, (2) training with a vision-only backbone, and (3) training with a vision-language backbone. Additionally, we indicate whether each method supports zero-shot (ZS) classification. This capability is marked as (✓) for CLIP-based approaches and (✗) for methods relying solely on vision backbones. We also denote the methods with learnable backbone (LB) by (✓) and the prompt- or adapter-based methods having frozen backbone by (✗) [^4] Fine-tuning results for the frozen backbones are reported in Appendix C.. Under the standard 10-step class-incremental split, our method achieves a final accuracy of 83.3%, outperforming the previous state-of-the-art by 1.3% points. Furthermore, for more challenging incremental sequences of 20 and 50 steps, our approach consistently surpasses RanPAC, the strongest existing baseline in terms of the ‘Last’ accuracy. It is worth noting that RanPAC operates in a high-dimensional space ($d=10000$), in contrast to most other methods which use default embedding dimensions for classification. Moving to larger-scale continual learning scenarios (Table III), our method also demonstrates strong performance. On the ImageNet_subset dataset, it achieves superior top-5 Last accuracy, surpassing the performance of linear probing [19], and yields comparable top-1 Last accuracy. On the more challenging ImageNet1000, our method outperforms the state-of-the-art approach by 1.2% in top-1 Last accuracy, while also achieving superior results in top-5 accuracy. Overall, our experimental results demonstrate that the proposed method consistently outperforms state-of-the-art approaches on all three datasets, showcasing its superior performance and stability in continual learning settings.
TABLE V: Last accuracy on ImageNet-A ($20\times 10$), ImageNet-R ($20\times 10$) and VTAB ($10\times 5$) .
Method RanPAC CODA-Prompt ConvPrompt EASE SPU CLAP4CLIP RAPF PROOF Ours
ImageNet-A 57.9 43.8 55.6 51.9 55.4 45.6 38.5 45.1 58.7
ImageNet-R 78.0 74.7 78.6 76.8 79.6 80.7 78.7 74.7 82.1
VTAB 90.2 78.9 61.9 89.2 93.0 91.5 81.7 72.5 94.0
@@
TABLE VI: The effect of different components. ‘FT’ represents only applying contrastive loss for training, ‘One-hot Label’ utilizes the prompt of each class as a classification head during training, and uses the one-hot encoded labels to calculate the cross-entropy loss. ‘Naive KD’ distills the output of the current data between old and new models with the KD loss proposed in[36] (ref Eq. 11).
@@
To further validate the effectiveness of the proposed method, we conduct experiments on eight fine-grained datasets. In fine-grained datasets, all categories belong to the same parent category in a hierarchical relationship, which intuitively indicates a high degree of semantic similarity among them, which is more challenging. In Table IV, we present the performance of our proposed method compared to other methods on different splits of the datasets. Our method achieves the best performance on five datasets, with improvements of up to 4.3%, and delivers competitive results on the remaining benchmarks. Moreover, our method even surpasses the results of linear probing evaluation on multiple datasets, which is typically used to assess the quality of pre-trained features.
The results of our method combined with the RanPAC classifier are presented in Appendix A, further validating the strength of the learned feature representations.
5.3 Generalization Performance with Semantic Gaps
We have conducted experiments on three challenging datasets known for their significant domain shift from pre-training data: ImageNet-A [67], ImageNet-R [68] and VTAB [72, 69]. ImageNet-A primarily consists of ‘natural adversarial examples’ or hard samples where models often struggle. ImageNet-R contains renditions of ImageNet classes in various artistic styles (e.g., paintings, sketches, cartoons), presenting a substantial stylistic shift. VTAB [69] was collected from 5 datasets with distinctly different distributions, including: the RESISC45 remote sensing image scene classification dataset, the DTD texture description dataset, the Pets dataset, the EuroSAT satellite imagery dataset, and the Flowers dataset. Each dataset is assigned to a task and contains 10 classes. These datasets are widely used in recent class-incremental learning research to evaluate model performance when faced with a significant domain gap between pre-train and downstream data [69, 25, 73].
The results of the experiments are in Table V. These results demonstrate that our method achieves competitive/superior performance on all datasets, indicating its robust generalization capability even when faced with considerable semantic and stylistic gaps between the pre-training and downstream data. Compared to all added baselines, our method shows better resilience to such domain shifts.
@@ 5.4 Few-shot Continual Learning Setting
Additionally, we evaluate our method alongside several baselines to assess their ability to preserve zero-shot performance in multi-domain task-incremental learning. Detailed results are presented in Appendix B.
@@ 5.5.1 Components of our proposed method
We conduct ablation experiments on CIFAR100 and ImageNet_subset to thoroughly examine the effect of the different proposed modules. Both datasets were split into 10 steps, and the results are shown in Table VI. On both datasets, the contrastive loss significantly improves the final accuracy compared to Continual-CLIP. However, when we additionally apply the one-hot encoded labels to add an auxiliary classification loss, we observe a significant performance drop. By incorporating our ‘SG-RL’ module, the model shows a gain of 1.8 points on CIFAR100 and 0.7 points on ImageNet_subset. Furthermore, when we apply knowledge distillation to the loss function, our proposed ‘SG-KD’ demonstrates superior performance over ‘Naive KD’ (ref Eq. 11) on both datasets, with a gain of 0.9 points on CIFAR100.
TABLE VII: Comparison of different vision models on ImageNet_subset, learned with 10 steps.
ImageNet_subset
Avg Last
ResNet 50 C-CLIP[26] - 65.6
Ours 80.2 71.3
ViT-B/16 C-CLIP[26] - 75.2
Ours 89.8 83.1
ViT-L/14 C-CLIP[26] - 81.5
Ours 91.9 85.8
TABLE VIII: Comparison of training different layers of image encoder.
@@
Figure 5: The comparison of accuracy when saving different number of exemplars on ImageNet_subset.
@@ 5.5.2 Validation on various vision models
Our primary experiments employed ViT-B/16 as the image encoder. However, it is important to investigate the effectiveness of our method with other architectures as well. Hence, we validate our approach on two additional vision models: ResNet 50 and ViT-L/14. Table VII presents the comparison between Continual-CLIP (illustrated as ‘C-CLIP’ for short in the table) evaluation and our proposed method on ImageNet_subset within 10 steps. Despite the differences in architecture and parameter count, all three image encoders exhibit significant improvements compared to Continual-CLIP evaluation. These results suggest that our method is effective across different scales of image encoders, irrespective of architecture.
TABLE IX: Ablation study on hyper-parameters sensitivity. (‘I’ stands for ‘ImageNet_subset’.)
@@ 5.5.3 Effect of training different layers of image encoder
From Table VIII, we observe that when training the last 6 blocks of the image encoder (half of the blocks in ViT-B/16), the model fails to overcome catastrophic forgetting. However, significant improvements are achieved when our proposed SG-RL and SG-KD modules are applied. Training only the linear projection layer yields decent results, but it appears that the model’s potential is greatly limited with such a small number of parameters. To explore the impact of incorporating more parameters into the training process, we gradually training additional blocks. It is evident that the best performance is achieved when only the last blocks are trained.
@@ 5.5.5 Hyper-parameters sensitivity
We conduct sensitivity analysis experiments on the two proposed losses to assess their impact. It can be seen from Table IX, our method shows strong robustness to both hyper-parameters ranged from 0.1 to 1. The best results are obtained when $\lambda_{1}$ and $\lambda_{2}$ are set to 0.5, 0.1 respectively.
@@
5.6 Comparison with one-hot label
@@
TABLE XI: Results compared to one-hot label on two general datasets with different task splits.
Dataset ImageNet_subset ImageNet1000
split $10\times 10$ $100\times 10$
One-hot Label 73.2 64.3
Ours 83.1 75.1
@@ 6 Conclusion
Appendix A Discussions with RanPAC
When combined with a RanPAC-style classifier, our model consistently outperforms the original RanPAC method, demonstrating the robustness and transferability of the learned features. Under an oracle setting—where RanPAC’s classifier is constructed using full data statistics—our approach achieves even higher performance, further indicating that the feature representations learned by our method possess strong linear separability.
TABLE XII: Combination with RanPAC-style classifier. $\dagger$ denotes RanPAC classifier adaptation using exemplars with continually updated feature extractor.
Dataset CIFAR100 Aircraft Cars Dogs ImageNet-R
#classes 100 100 196 120 200
split $10\times 10$ $10\times 10$ $28\times 7$ $12\times 10$ $20\times 10$
RanPAC [46] 82.0 62.3 86.3 77.0 78.0
Ours $+$ RanPAC [46] 84.8 69.3 88.9 78.5 82.4
Ours $+$ Oracle RanPAC [46] 86.6 71.4 91.9 81.5 83.8
Appendix B Evaluation on Multi-Domain Task Incremental Learning Setting
The Multi-Domain Task Incremental Learning (MTIL) setting was first introduced by ZSCL[78], which proposed a knowledge distillation module and a weight ensemble mechanism to mitigate zero-shot performance degradation. MTIL focuses on continual learning across different datasets, with an emphasis on preserving zero-shot capabilities. In contrast, our class-incremental learning setting aligns with prior works, aiming to continually learn new tasks while retaining knowledge of previously learned classes.
We compared our method with these methods mentioned above in Table XIII and Table XIV following the same datasets order I and order II in ZSCL. The backbones of all methods are CLIP ViT-B/16 for fair comparison. Under MTIL setting, methods are evaluated with Transfer, Last and Average three main metrics, corresponding to zero-shot performance, averaged accuracy after last task and averaged incremental accuracy respectively. As shown in Table XIII, our method achieves optimal or second-best average performance across all evaluation metrics. Notably, it yields a 1.2% improvement on the Average metric, while being only 0.3% and 0.2% lower than TreeProbe on the Last and Transfer metrics, respectively. The results for Order II are presented in Table XIV, where our method consistently achieves superior average performance across all evaluation metrics.
TABLE XIII: Comparison with different methods on the MTIL benchmark (Order I) in terms of ‘Transfer’, ‘Last’, and ‘Average’ performance(%).
Methods
Aircraft
Caltech101
CIFAR100
DTD
EuroSAT
Flowers
Food
MNIST
OxfordPets
Cars
SUN397
Average
Transfer ZSCL [78] - 86.0 67.4 45.4 50.4 69.1 87.6 61.8 86.8 60.1 66.8 68.1
MoE4CLIP [79] - 87.9 68.2 44.4 49.9 70.7 88.7 59.7 89.1 64.5 65.5 68.9
TreeProbe [80] - 87.9 68.2 45.3 54.6 71.1 88.9 59.5 89.1 64.6 64.1 69.3
Ours - 84.3 68.0 46.1 56.0 71.3 87.0 57.9 91.0 64.6 65.0 69.1
Last ZSCL [78] 40.6 92.2 81.3 70.5 94.8 90.5 91.9 98.7 93.9 85.3 80.2 83.6
MoE4CLIP [79] 49.8 92.2 86.1 78.1 95.7 94.3 89.5 98.1 89.9 81.6 80.0 85.0
TreeProbe [80] 52.5 95.6 81.9 66.9 95.6 95.6 92.2 98.6 93.2 86.2 81.8 85.5
Ours 64.5 93.6 79.8 70.8 93.2 94.9 91.1 98.0 92.9 81.1 77.6 85.2
Average ZSCL [78] 45.1 92.0 80.1 64.3 79.5 81.6 89.6 75.2 88.9 64.7 68.0 75.4
MoE4CLIP [79] 50.2 91.9 83.1 69.4 78.9 84.0 89.1 73.7 89.3 67.7 66.9 76.7
TreeProbe [80] 52.5 93.6 79.5 61.9 80.5 78.0 90.4 73.7 90.2 68.5 65.7 75.9
Ours 66.6 93.3 79.3 65.8 80.3 84.2 89.2 72.6 91.8 67.9 66.2 77.9
Appendix C Results with Learnable Backbone
In our main experimental tableII, methods are distinguished by whether they employ a learnable backbone (LB), indicating if pre-trained weights are updated during continual learning. A significant portion of existing state-of-the-art methods operate with a frozen backbone. Our proposed approach, in contrast, involves fine-tuning part of the pre-trained weights of the Vision Transformer. This naturally raises a question regarding the source of our observed performance gains: to what extent are they attributable to our novel semantic guidance techniques (i.e., SG-KD and SG-RL) versus the general benefits derived from fine-tuning a segment of the backbone? To address this and provide a more direct comparison of the core methodological contributions, we have included an additional comparative experiment in TableXV. In this supplementary analysis, we adapted several leading competitor methods, which originally used a frozen backbone, to also fine-tune the last residual block of their ViT backbone, mirroring the parameter update strategy of our method. While these adapted competitor methods predictably showed performance improvements over their frozen-backbone counterparts due to the increased learnability, our proposed method consistently maintained a significant advantage even under this condition. This finding strongly suggests that the primary driver of our method’s superior performance stems from the effective exploitation of semantic knowledge via our proposed SG-KD and SG-RL mechanisms, rather than being solely an artifact of partial backbone fine-tuning. This isolates and underscores the distinct benefits contributed by our novel techniques.
Appendix D More Details
D.1 Unified Exemplar Management for Replay
Among these compared methods, some of them (L2P, DualPrompt, RanPAC, CODA-Prompt, ConvPrompt, EASE, and RAPF) were originally proposed as exemplar-free CIL algorithms, some were exemplar-based CIL methods (SPU, CLAP, PROOF). CoOp and CoCoOp were proposed for prompt learning in a non-continual manner. Except EASE and RAPF, for all of these compared methods, we employed the herding algorithm to select and store 20 exemplars per old class. [^5] EASE employs prototype synthesis strategy to construct classifier, RAPF employs a decomposed parameter fusion strategy for its adapter module, these two methods are not well compatible with exemplars replay. These exemplars were then replayed during the training of subsequent incremental tasks, and their loss was incorporated into the total training objective for each method. For CoOp and CoCoOp, we adapted them to the CIL setting and also applied the same exemplar replay strategy.
D.2 Hyperparameter and Training Parameter Settings
For each baseline method, we prioritized using the optimal hyperparameters and training configurations reported in their respective original publications whenever possible. If the original parameters were not directly applicable to our specific CIL setup or datasets, or if complete details were not provided, we conducted a reasonable hyperparameter search (e.g., on a validation set) to identify settings that yield the best performance for each method under our evaluation framework.
By adhering to these careful implementation practices and standardized evaluation protocols, we aimed to establish a robust and equitable comparison platform for all methods, ensuring that the observed performance differences accurately reflect their capabilities in the CIL context.
TABLE XIV: Comparison with different methods on the MTIL benchmark (Order II) in terms of ‘Transfer’, ‘Last’, and ‘Average’ performance(%). Results of TreeProbe on order II are reproduced with their official published code, as this experiment was not reported in the original paper.
Methods
Cars
Food
MNIST
OxfordPets
Flowers
SUN397
Aircraft
Caltech101
DTD
EuroSAT
CIFAR100
Average
Transfer ZSCL [78] - 88.3 57.5 84.7 68.1 64.8 21.1 88.2 45.3 55.2 68.2 64.1
MoE4CLIP [79] - 88.8 59.5 89.1 69.9 64.6 18.1 86.9 43.7 54.6 68.2 64.3
TreeProbe* [80] - 87.9 59.5 89.1 71.1 61.1 24.7 87.2 43.5 54.7 62.7 64.2
Ours - 87.0 58.1 90.8 71.4 65.1 23.8 88.5 46.4 55.1 68.3 65.5
Last ZSCL [78] 78.2 91.1 97.6 92.5 87.4 78.2 45.0 92.3 72.7 96.2 86.3 83.4
MoE4CLIP [79] 84.1 88.5 94.0 91.8 94.1 77.8 50.4 93.3 77.1 87.7 86.6 84.2
TreeProbe* [80] 86.3 92.2 98.5 91.5 81.0 79.0 49.7 90.2 62.7 88.2 81.9 81.9
Ours 86.6 90.6 96.2 93.1 94.9 74.0 57.2 93.2 69.4 94.7 82.2 84.7
Average ZSCL [78] 81.7 91.3 91.1 91.0 82.9 72.5 33.6 89.7 53.3 62.8 69.9 74.5
MoE4CLIP [79] 84.9 89.9 89.3 91.4 86.2 72.2 33.4 89.4 53.3 61.4 69.9 74.7
TreeProbe* [80] 86.3 91.8 91.4 91.1 83.1 71.7 36.0 89.2 49.8 60.9 64.5 74.3
Ours 88.2 91.0 90.4 93.0 86.4 71.1 39.7 90.5 53.6 62.6 69.6 76.0
TABLE XV: Results with a learnable backbone across multiple evaluated datasets.
Dataset LB CIFAR100 ImageNet_sub ImageNet1000 Food Cars Aircraft Pets Caltech Flowers CUB Dogs
#classes 100 100 1000 101 196 100 37 102 102 200 120
CoOp[55] 68.1 70.2 66.7 84.6 86.1 57.6 94.0 90.2 97.0 80.1 75.8
CoCoOp[54] 65.9 71.0 56.0 79.1 84.9 59.8 92.8 88.2 97.2 77.0 69.7
L2P[14] 65.9 72.7 63.3 81.7 86.9 61.1 92.8 88.6 97.0 82.6 75.9
DualPrompt[15] 65.1 68.4 64.0 80.5 82.7 45.5 90.6 86.4 96.7 79.6 70.3
Ours 83.3 83.1 75.1 90.9 88.6 66.6 94.0 92.3 96.2 83.7 79.6
  • CoOp We adopted the class-specific prompts variant due to its superior performance. The number of context words (i.e., the number of learnable prompt tokens) was set to 16 per class. The learning rate was set to 0.01, batch size was 256, and each task was trained for 10 epochs.
  • CoCoOp We set the number of prompt tokens to 16. The meta-network consisted of two linear layers with a ReLU activation function between them. The learning rate was 0.001, the batch size was 8, and training spanned 10 epochs per task.
  • L2P We set the prompt length to 5 and prompt pool size equal to the number of tasks, meaning each task had a corresponding set of prompts. The trade-off for loss between queries and learnable keys was 0.1. A learning rate of 0.03 and a batch size of 16 were used. Standard datasets were trained for 5 epochs per task, extended to 10 epochs for fine-grained datasets to ensure convergence.
  • DualPrompt We set the length of general prompts to 5 and applied them to the first two transformer layers, while expert prompts, also of length 5, were used in the subsequent three layers. The prompt pool size was set equal to the number of tasks. The learning rate was 0.03 with a batch size of 24. General datasets were trained for 5 epochs per task, extended to 10 epochs for fine-g rained datasets.
  • RanPAC We set $M$ to 10,000, which is utilized to construct the ramdom matrix for mapping features to high-dimensional space.
  • CODA-Prompt We set the length of prompt to 8, the prompt pool size to match the number of classe, the feature dimension of the prompts is 768. We set the learning rate to 0.001 and the batch size to 128, training for 20 epochs per task. We used the Adam optimizer with a cosine learning rate scheduler.
  • ConvPrompt We configured a kernel size of 17. It learned 5 prompts per task, each of length 20, and the prompt pool size was equal to the number of tasks. Only expert prompts were utilized, applied in the first 7 transformer layers.
  • EASE For CIFAR100, the learning rate was 0.025, batch size 48, and weight decay 0.0005. Other datasets adopted the ImageNet-R configuration: learning rate 0.05, batch size 16, and weight decay 0.005. All datasets were trained for 20 epochs per task using the SGD optimizer with a cosine annealing scheduler. The adapter’s intermediate dimension was 64, and the ensembling coefficient $\alpha$ between different subspaces was 0.1.
  • SPU We set the selection rate to 0.1 and the percentage of samples per task used for calculating parameter importance to 0.25. The learning rate was 7.5e-6, the batch size was 64, and the weight decay was 0.2. We employed the AdamW optimizer with a cosine annealing scheduler.
  • CLAP the first stage training used a learning rate of 0.001, a batch size of 32, and the SGD optimizer with a cosine annealing scheduler for 5 epochs per task. The second stage (class-balanced consolidation) trained for 2 epochs with a learning rate of 0.0001. The number of Monte Carlo forward passes sampled from probabilistic adapters was 20. A single-layer transformer-decoder was used in the VGA Module, and task-specific probabilistic adapters were learned for each task. The coefficient $\lambda$ for the prior-matching DKL term was 0.001, and $\gamma$ for the past-task distribution regularization loss was 15.
  • PROOF We use single linear layer for the task-specific MLP for each task. The initial learning rate was set to 0.001, the batch size to 64, and the weight decay to 0.05. We utilized the SGD optimizer with a cosine annealing scheduler.
For methods based solely on a vision encoder (L2P, DualPrompt, RanPAC, CODA-Prompt, ConvPrompt, and EASE), we utilized the vision features directly for classification. This involved discarding the projection layer from the pre-trained CLIP weights that is originally used to map vision features to the language space. Conversely, for methods designed for or leveraging vision-language models (SPU, CLAP, RAPF, and PROOF, as well as our adaptations of CoOp and CoCoOp), we employed all pre-trained weights of the CLIP model.